### v2

# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

Abstract
Algorithmic scatter dimension is a notion of metric spaces, introduced recently by Abbasi et al. (FOCS 2023), which unifies various well-known metric spaces, including continuous Euclidean space, bounded doubling space, planar and bounded treewidth metrics. Recently, Bourneuf and Pilipczuk (SODA 2025) showed that metrics induced by graphs from any fixed proper minor closed graph class have bounded scatter dimension. Abbasi et al. presented a unified approach to obtain EPASes (i.e., $(1+\varepsilon)$-approximations running in time FPT in $k$ and $\varepsilon$) for $k$-Clustering in metrics of bounded scatter dimension. However, a seemingly inherent limitation of their approach was that it could only handle clustering objectives where each point was assigned to the closest chosen center, i.e., Voronoi clustering. They explicitly asked, if there exist EPASes for constrained $k$-Clustering for metrics of bounded scatter dimension.
We present a unified framework which yields EPASes for metrics of bounded algorithmic scatter dimension, for capacitated and fair $k$-Median/Means. Our framework exploits coresets for such constrained clustering problems in a novel manner, and notably, requires only coresets of size $\left(k\log n\varepsilon^{-1}\right)^{O(1)}$, which are usually constuctible even in general metrics. Furthermore, we note that due to existing lower bounds it is impossible to obtain such an EPAS for Capacitated $k$-Center, thereby, essentially answering the complete spectrum of the question.
Our results on capacitated and fair $k$-Median/Means make significant progress, and provide the first EPASes for these problems in broad families of metric spaces. Earlier such results were only known in continuous Euclidean spaces due to the work of Cohen-Addad & Li, (ICALP 2019), and Bandyapadhyay, Fomin & Simonov, (ICALP 2021; JCSS 2024), respectively. Along the way, we obtain faster EPASes for uncapacitated $k$-Median/Means, improving upon the running time of the algorithm by Abbasi et al. (FOCS 2023).
1 Introduction
Clustering plays an indispensable role in our understanding of very large scale data. In $k$-clustering problems, we are given a set $\mathcal{P}$ of data points in a metric space $\mathcal{M}$, and the goal is to partition $\mathcal{P}$ into $k$ partitions (namely, clusters) such that nearby points should ideally be in the same part, and each cluster is represented by a center. Clustering problems have been studied extensively in theoretical computer science, machine learning, computer vision, among other fields. Among the various notions of clustering, $k$-Median and $k$-Means are arguably the most popular and have been studied algorithmically for decades [KMN+02, HM04a, BCP+24, CEMN22, HM04b, BHI02, KSS10]. Clustering under capacity constraints is a natural and fundamental problem, and hence, quite conceivably, Capacitated $k$-Median and Capacitated $k$-means have garnered lot of attention in recent years; see [BFRS15, CR05, Li17, Li16, DL16, CGTS02, Coh20].
Capacitated $k$-Median and Capacitated $k$-means are inherently harder than the vanilla (Uncapacitated) $k$-Means or $k$-Median, since there is hard constraint on the number of clusters as well as the number of points that can be assigned to each cluster. In fact, till-date the best known approximation for capacitated $k$-median remains $O(\log k)$ (folklore), using the probabilistic embedding of metric spaces into trees (Adamczyk et al. gave an explicit proof in [ABM+19] for completeness). Constant-factor approximation algorithms exist when either the capacities [Li17, Li16], or the number of clusters [DL16, BRU16], are allowed to be exceeded by a $(1+\varepsilon)$ factor, for some constant $\varepsilon>0$. From the lower bounds side, surprisingly, no better lower bound than $(1+\frac{2}{e}-\varepsilon)$ is known, which is inherited from the uncapacited version. Given this evident gap, it is natural to ask whether the approximation landscape can be improved for metric spaces with special properties, possibly at the cost of higher running time. While it remains an open question whether a constant-factor approximation algorithm exists for general metrics that runs in polynomial time, Cohen-Addad and Li [CL19a], in 2019, showed a $(3+\varepsilon)$-approximation for the problem, running in time FPT in $k$ and $\varepsilon$. However, even in the uncapacitated case and with FPT running time, the lower bound remains at $1+\frac{2}{e}$, assuming Gap-ETH. Notably, for uncapacitated case, this lower bound can be matched by an FPT algorithm [CGK+19]. A central open question, thus, is: For which families of metric spaces, is it possible to obtain better upper bounds for capacitated $k$-Median/Means.The only improvement in the affirmative sense is known for Euclidean spaces. For instance, for Euclidean spaces, Cohen-Addad and Li showed that there exists $(1+\varepsilon)$-approximation algorithms for both problems with running time $f(\varepsilon,k)n^{O(1)}$ [CL19a] (see also [BJK18]).  [^1] In this context, a recent work of [HLLW25] on coresets for constrained clustering is also relevant. We discuss the implications of this work and its relation with our work in Section 4.4.
As mentioned before, stronger upper bounds are known for uncapacitated $k$-Median/Means. Indeed, Cohen-Addad et al. [CGK+19] obtained tight $(1+\frac{2}{e}+\varepsilon)$ and $(1+\frac{8}{e}+\varepsilon)$ approximation factors for $k$-Median and $k$-Means in general metric spaces, respectively, with running times FPT in $k$ and $\varepsilon$. However, an important question remains: Can the lower bound barriers for uncapacitated $k$-Median/Means be broken for special metric spaces? To that end, improved FPT-approximation algorithms have been developed for Euclidean spaces; see [BHI02, KSS10]. Bădoiu, Har-Peled and Indyk [BHI02] showed that, for points in $\mathbb{R}^{d}$, a $(1+\varepsilon)$-approximate $k$-Median can be computed in expected $2^{{k/\varepsilon}^{O(1)}}d^{O(1)}n\log^{O(k)}n$ time. Subsequently, Har-Peled and Mazumdar in their major work [HM05], showed the existence of small coresets for the problems of computing $k$-Median and $k$-Means clustering for points in low dimension. Kumar, Sabharwal and Sen in their breakthrough work [KSS10] gave simple randomized algorithms for $k$-Median/Means that obtain $(1+\varepsilon)$-approximation solutions with constant probability, running in time $O(2^{{(k/\varepsilon)}^{O(1)}}dn)$. A fundamental aspect of these works is that they primarily rely on key sampling ideas (or, roughly speaking, coresets) that crucially exploits properties of the Euclidean space. Moreover, all these works primarily focus on the continuous Euclidean cases (where centers can be opened anywhere), except that the result of Kumar et al. [KSS10] also works for discrete $k$-Means. Feldman et al. [FSS20] obtained coresets for $k$-Means, PCA and projective clustering whose sizes are independent of the dimension. Later, Sohler and Woodruff [SW18] generalized these results and obtained coresets for $k$-Median and for subspace approximation of sizes independent of the dimension $d$. Braverman et al. [BJKW21] designed coresets for for $k$-Median in (the shortest-path metric of) an excluded-minor graph, which is also applicable to Euclidean metrics by using terminal embedding result of Narayanan and Nelson [NN19] that extends the Johnson-Lindenstrauss Lemma. After a long sequence of work [CAFS21, HV20, BBH+20, BJKW21, FL11], finally in 2021, Cohen-addad, Saulpic, and Schwiegelshohn [CSS21] gave a unified framwork for a large variety of metrics, ranging from Euclidean space, doubling metric, minor-free metric, as well as capturing the general metric case. However, while these results provide excellent bounds on coreset size, a key concern remains: the computational cost of clustering using them. Hence, a natural question to ask: Can coresets be used to obtain faster algorithms for $k$-Clustering?
EPASes for (Constrained) Clustering.
Efficient parameterized approximation scheme (EPAS) [^2] A $(1+\varepsilon)$ approximation algorithm that runs in time $f(k,\varepsilon)\operatorname{poly}(n)$ for every $\varepsilon>0$. is perhaps, the best combination of approximation and parameterized algorithms, and has been a majorly popular theme in recent algorithms research (see [ABB+23, BFM24, Mat00, ORSS13, DX20] and the references therein). However, for $k$-Median/Means, unfortunately, it is not possible to obtain EPASes for points in general metric spaces [Das08, ACKS15]. Therefore, quite conceivably, much of the work focused on designing EPASes for points in special metrics, e.g., continuous Euclidean spaces [KSS10, BHI02]. In a recent work [ABB+23], Abbasi et al. presented a unified EPAS which works for various clustering objectives as well as diverse metric spaces. To that end, they introduced the notion of $\varepsilon$-scatter dimension (see Section 2.1 for definition), which includes various commonly studied metric spaces, including, continuous Euclidean space, doubling spaces, etc. Very recently, Bourneuf and Pilipczuk [BP25] showed that metrics induced by graphs from any fixed proper minor-closed graph class also have bounded scatter dimension. Moreover, they introduced metric analogs of well-known graph invariants from the theory of sparsity, and as a consequence of this toolbox showed a coreset for $k$-Center in any proper minor-closed graph class whose size is polynomial in $k$ but the exponent of the polynomial depends on the graph class and $\frac{1}{\varepsilon}$. Abbasi et al. [ABB+23] bypassed coresets to obtain their clustering results. They mentioned that, obtaining clustering algorithms via coresets faces information-theoretic limitation, particularly in the context of $k$-Center, since coresets of desirable sizes do not exist in high-dimensional Euclidean spaces [BJKW19]. However, a seemingly inherent limitation of their approach was that they could only handle clustering objectives, where each point was assigned to the closest chosen center, i.e., Voronoi clustering [^3] A recent work of Gadekar and Inamdar [GI25] also used scatter dimension to design an EPAS for a common generalization of $k$-Center and $k$-Median that is not captured by the general norm framework of [ABB+23]; demonstrating that the notion of (algorithmic) scatter dimension has a wider applicability in clustering. However, the clustering objective considered in [GI25] still can only handle Voronoi assignments.. Indeed, in their work they explicitly asked the following open question:
Question 1.
Is it possible to design EPASes for constrained $(k,z)$-Clustering in metric spaces of bounded (algorithmic) scatter dimension, for assignment constraints such as capacities, fairness, and diversity?
In fact, [ABB+23] asked this question for a more general problem, called constrained $k$-Clustering. However, this problem is so general that it also captures Capacitated $k$-Center, which is known to be W$[1]$-hard to approximate to a factor arbitrarily close to $1$ in metric spaces of bounded highway dimension [FV22]. Furthermore, it has been observed that metric spaces of bounded highway dimension have bounded scatter dimension [^4] See the talk on [ABB+23] at FOCS 2023: https://focs.computer.org/2023/schedule/., implying that the original question of [ABB+23] cannot be answered in the affirmative. Therefore, we reinterpret their question by restricting our focus to constrained $(k,z)$-clustering, where the objective is to minimize the sum of $z$th power of distances. This problem is general enough to capture many well studied clustering problems (e.g., $z=1,2$, correspond to $k$-Median and $k$-Means, respectively), but also manages to bypass the hardness of result of [FV22].
1.1 Our Contributions
A Unified Framework for Constrained Clustering.
In this work, we answer 1 in the affirmative by designing a unified framework that yields EPASes for a large class of constrained $(k,z)$-clustering in metric spaces with bounded algorithmic scatter dimension. Somewhat informally, we require that the constrained $(k,z)$-clustering problem satisfies the following requirements.
  1. 1.
    There exists an algorithm $\mathcal{A}_{C}$ that computes a coreset $(Y,w)$ for the problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$.  [^5] We emphasize here that, the required bound of $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ on the coreset size is attainable even in general metric spaces for many natural constrained $(k,z)$-clustering problems; thus we do not need to rely on metric-specific properties to obtain coresets whose size independent of $\log n$.
  2. 2.
    There exists an algorithm $\mathcal{A}_{A}$ that takes as an input any weighted set of points $Z$ and candidate set of $k$ centers, $C$. Either it finds a near-optimal assignment out of all feasible (i.e., respecting all problem-specific constraints) assignments from $Z$ to $C$, or outputs that no feasible assignment from $Z$ to $C$ exists.
  3. 3.
    There exists an algorithm $\mathcal{A}_{B}$ that takes as an input a finite set of requests $Q=\{(p,r):p\in Y,r\in\mathbb{R}\}$, and an $\varepsilon\in(0,1)$, and either returns a center $c\in F$ such that (i) ${\sf d}(p,c)\leq(1+\varepsilon)r$ for all requests $(p,r)\in Q$, and (ii) $c$ satisfies problem-specific feasibility requirements; or correctly outputs that no such center exists.
We call such problem as Well Constrained $(k,z)$-Clustering problem. We allow each of the three algorithms to run in time FPT in $k,\varepsilon$, and potentially some problem-specific parameters. As we shall see later, these requirements are extremely mild and are satisfied by many natural clustering problems. Our main result is stated in the following theorem.
Theorem 2 (Informal version of Theorem 10).
There exists a $(1+\varepsilon)$-approximation for Well Constrained $(k,z)$-Clustering running in time $f(k,\varepsilon)\cdot|\mathcal{I}|^{O(1)}$ in metric spaces of bounded algorithmic scatter dimension, where $|\mathcal{I}|$ denotes the size of the input instance $\mathcal{I}$. In addition, if the coresets for the problem can be computed deterministically, then the algorithm is deterministic.
Now, we briefly discuss some of the consequences of our framework. A summary of these results is given in Table 1.
Constraint FPT Factor Property Theorem Section
Capacities uniform/non-uniform $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))})$ Deterministic* Thm. 3 4.1
$(\alpha,\beta)$-Fairness Also: chromatic & $\ell$-diversity $\exp{(O(\frac{k\Gamma\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Gamma\Lambda}{\varepsilon}))})$ Deterministic* Thm. 4 4.2
Vanilla $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k}{\varepsilon}))})$ Randomized Thm. 5 4.3
Table 1: Consequences of our EPAS framework for Well Constrained $(k,z)$-Clustering. In the second column, only the FPT part of the running time is specified, and $\Lambda\coloneqq\lambda(\varepsilon/c)$, where $\lambda(\cdot)$ is the (algorithmic) scatter dimension of the metric space, and $c$ is a constant only depends on $z$. In the second row, $\Gamma$ represents the number of equivalence classes w.r.t.  colors (see Definition 15). For the results marked as “Deterministic*”, if the coresets for the problem can be computed deterministically, then our EPAS is completely deterministic. Note that the results of [ABB+23] only work for the vanilla version, and the FPT factor of their randomized EPAS is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))$; thus our EPAS shaves off a $\log k$ factor from the exponent.
EPASes for Capacitated and Fair $(k,z)$-Clustering:
As a consequence of our framework, we obtain the first EPASes for Capacitated $(k,z)$-Clustering in metric spaces of bounded algorithmic scatter dimension. Prior to this, EPASes for Capacitated $k$-Median/Means were known only in Euclidean Spaces [CL19b, BJK18]. Furthermore, our results extend to the fair version of the problems (namely, $(\alpha,\beta)$-fair clustering [^6] In this version, points are divided into multiple (potentially overlapping) groups, and we are additionally given upper and lower bounds on the number of points of each group that can belong to each of the $k$ clusters. The goal is to find a minimum-cost clustering of points satisfying these fairness constraints. A formal definition is given in Section 4.2.  [BFS24, BGJ21]) In what follows, we state our theorems, and the key ingredients and proofs are given in Section 4.1 and Section 4.2 respectively.
Theorem 3.
For any fixed $z\geq 1$, Capacitated $(k,z)$-Clustering admits an EPAS in metric spaces of algorithmic $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$ running in time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$, where $\Lambda\coloneqq\lambda({\varepsilon}/{c})$, and $c$ is a constant that only depends on $z$.
Theorem 4.
$(\alpha,\beta)$-Fair $(k,z)$-Clustering admits an EPAS in metric spaces of algorithmic $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$ running in time $\exp{(O(\frac{k\Gamma\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Gamma\Lambda}{\varepsilon}))})\cdot|\mathcal{I}|^{O(1)}$, where $\Lambda\coloneqq\lambda({\varepsilon}/{c})$, and $c$ is a constant that only depends on $z$.
Faster EPASes for Vanilla $(k,z)$-Clustering.
Since Vanilla (a.k.a. uncapacitated/Voronoi) $(k,z)$-Clustering is a special case of Capacitated $(k,z)$-Clustering, Theorem 3 implies an EPAS for this variant with running time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$. We note that the running time of the EPAS from [ABB+23] is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))\cdot|\mathcal{I}|^{O(1)}$. Thus, note that our running time is already better than that of [ABB+23] by a factor of $\log k$ in the exponent, but we incur an additional $\log\Lambda$ factor. However, in Section 4.3, we describe a simple modification to our framework, where we replace a key deterministic branching step by a randomized one, leading to an improvement in the running time. Specifically, compared to the EPAS of [ABB+23], our resulting running time matches the dependence on $\varepsilon$, and shaves off a $\log k$ factor from the exponent. More formally, we prove the following theorem in Section 4.3.
Theorem 5.
For any fixed $z\geq 1$, $(k,z)$-Clustering admits an EPAS in metric spaces of algorithmic $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$ running in time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$, where $\Lambda\coloneqq\lambda({\varepsilon}/{c})$, and $c$ is a constant that only depends on $z$.
EPASes for bounded highway dimension.
We mention the following interesting corollary of our framework.
Corollary 6.
For a fixed $z\geq 1$, Well Constrained $(k,z)$-Clustering admits an EPAS in metrics of bounded highway dimension.
We find this corollary interesting due to the following reason. As mentioned earlier, metrics of bounded highway dimension are contained in the class of metrics of bounded scatter dimension. Therefore, our framework implies EPASes for Well Constrained $(k,z)$-Clustering for metric spaces of bounded highway dimension. Specifically, this includes Capacitated $(k,z)$-Clustering, which is in stark contrast with Capacitated $k$-Center, for which the result of [FV22] rules out such an EPAS (or even a PAS, i.e., an algorithm running in time $f(k)\cdot|\mathcal{I}|^{g(\varepsilon)}$).
1.2 Technical Overview
We briefly describe the technical ideas used to obtain the result, focusing mainly on constrained $k$-Median, for simplicity. Firstly, we compute a coreset $(Y,w)$ for the problem using algorithm $\mathcal{A}_{C}$, and focus our attention on obtaining a near-optimal solution on the coreset (which, in turn, implies a near-optimal solution for the original point-set). Let $\mathcal{C}=\{\mathcal{C}_{1},\dots\mathcal{C}_{k}\}$ be an optimal clustering of the coreset along with the corresponding centers $O=\{o_{1},\dots,o_{k}\}$, and let ${\sf OPT}$ be the optimal cost. We assume as have a good approximation of ${\sf OPT}$, which can be guessed approximately, up to factor $(1+\varepsilon)$, in polynomial time. On a high level, our framework maintains a a collection of $k$ request sets $(Q_{1},\dots,Q_{k})$ and a center $X=(x_{1},\dots,x_{k})$ satisfying the request sets. More specifically, each request set $Q_{i}$ is a collection of pairs of points and radii, $(p,\alpha)\in Y\times\mathbb{R}$, representing the demand that $p$ wants a center within distance $\alpha$ from itself. Additionally, the framework maintains an invariant that for $(p,\alpha)\in Q_{i}$, it must be that $p\in\mathcal{C}_{i}$, i.e., $p$ belongs to $i^{th}$ cluster in the optimal clustering $\mathcal{C}$. The main idea is to add requests to these request sets to force the algorithm refine its solution $X$ until it becomes a $(1+\varepsilon)$-approximate solution. In more detail, our framework can be conceptually broken down into multiple phases. The first phase is the leader guessing phase, which is the usual starting point in obtaining FPT-approximations for many clustering problems [CGK+19, CL19a]. In this phase, for each optimal center $o_{i}\in O$, we guess a point $p_{i}\in\mathcal{C}_{i}$ and its approximate distance [^7] Approximate up to a factor of $(1+\varepsilon)$. ${\sf d}(p_{i},o_{i})$, such that $p_{i}$ is closest to $o_{i}$ (ties broken arbitrarily). Since the size of the coreset is bounded, this guessing amounts to FPT many branches. Furthermore, this gives a baseline of minimum distances in each of the $k$ clusters in $\mathcal{C}$ that will be used by the algorithm later. Next, we add requests $(p_{i},d(p_{i},o_{i}))$ to $Q_{i}$ for all $i\in[k]$, and using algorithm $\mathcal{A}_{B}$, we compute a feasible set of centers $X$ satisfying all $Q_{i}$s. This means that, for each $i\in[k]$, center $x_{i}\in X$ satisfies ${\sf d}(p_{i},x_{i})\leq(1+\varepsilon){\sf d}(p,o_{i})$. Finally, we use algorithm $\mathcal{A}_{A}$ to compute a near-optimal assignment of points in $Y$ to $X$, to obtain a corresponding clustering $\mathcal{C}^{\prime}=(\mathcal{C}^{\prime}_{1},\dots,\mathcal{C}^{\prime}_{k})$. Then, as long as the cost of the current clustering $\mathcal{C}^{\prime}$ is larger than $(1+\varepsilon)\cdot{\sf OPT}$, we argue that there must exist an unhappy pair $(p,i)$, where $p\in\mathcal{C}_{i}$, i.e., $p$ is assigned to $o_{i}$ in the optimal solution, such that ${\sf d}(p,x_{i})>(1+\varepsilon/3){\sf d}(p,o_{i})$. Since, the size of coreset $Y$ is small, we can guess such unhappy pair $(p,i)$ using at most $|Y|\cdot k$ many branches. Next, we add the request $(p,\frac{{\sf d}(p,x_{i})}{1+\varepsilon/3})$ to request set $Q_{i}$. Finally, we recompute $X$ using $\mathcal{A}_{B}$ to additionally satisfy the new request, and repeat the whole process, if the cost of the recomputed $X$ is still larger than $(1+\varepsilon){\sf OPT}$. By carefully analyzing the nature of such unhappy pairs, we show the crucial property that for each request set $Q_{i}$, the radii of requests in $Q_{i}$ lie in an interval of width bounded by $O(1/\varepsilon)$. This, in turn, allows us to bound the number of requests that can be added to each $Q_{i}$ in metric spaces with bounded algorithmic scatter dimension. Thus, this leads to a recursive algorithm with branching factor $|Y|\cdot k$ and depth $O(\frac{k}{\varepsilon}\log(\frac{1}{\varepsilon})\cdot\lambda(\frac{\varepsilon}{c}))$, where $\lambda(\cdot)$ is the algorithmic scatter dimension of the metric space, and $c$ is a constant. Since $|Y|=\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$, this leads to the claimed running time by employing the standard trick, namely, $(\log n)^{t}=t^{O(t)}\cdot n^{O(1)}$.
Comparison to existing frameworks.
On a high level, our algorithm follows the outline of the framework recently proposed by [ABB+23], that unifies existing EPASes for Voronoi clustering problems. However, we face multiple challenges when we try to extend their framework to our problems. For instance, [ABB+23] introduce a notion of upper bounds, and devise them for their problems, that help bound the aspect ratio of radii in $Q_{i}$s. However, these upper bounds are inherently based on Voronoi properties, and it is not clear how to design them for our problems. This is where, we exploit the existence of general metric coresets. As mentioned before, coresets allows us to get a handle on the leaders and their distances, which in turn, allows us to bound the aspect ratio of radii in each request set, the task that upper bounds achieved in [ABB+23]. Another bottleneck that is apparent when extending their framework to our problems, is finding unhappy pairs. As mentioned before, if the cost of current solution $(X,\mathcal{C}^{\prime})$ is larger than $(1+\varepsilon){\sf OPT}$, there exists a unhappy pair, and we again exploit the compact size of coresets to find one. This is not a bottleneck for Voronoi clustering problems, since it can be shown that, in this case, there is a unhappy point $p\in Y$ such that ${\sf d}(p,X)>(1+\varepsilon/3){\sf d}(p,O)$. [^8] For a given center set $S$ and a point $p\in Y$, ${\sf d}(p,S):=\min_{s\in S}{\sf d}(p,s)$. This means that, if $p$ was assigned to $o_{i}$ in the optimal solution, then ${\sf d}(p,x_{i})\geq d(p,X)>(1+\varepsilon/3){\sf d}(p,o_{i})$, and hence $(p,i)$ is an unhappy pair, as required. Therefore, we can add a new request $(p,\tfrac{{\sf d}(p,x_{i})}{1+\varepsilon})$ to $Q_{i}$, so that the algorithm refines $x_{i}$ to satisfy the new request. However, for constrained clustering, it is possible that $p$ is assigned to $x_{j}\in X$ in $\mathcal{C}^{\prime}$, while it is assigned to $o_{i}$ in $\mathcal{C}$, for $i\neq j$, and it holds that ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$ due to the cost of $(X,\mathcal{C}^{\prime})$. Furthermore, it can happen that ${\sf d}(p,x_{i})$ is much smaller than ${\sf d}(p,x_{j})$ (the worst case is when ${\sf d}(p,x_{i})=0$, while ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$), which means that adding a request $(p,\tfrac{{\sf d}(p,x_{i})}{1+\varepsilon})$ does not refine $x_{i}$, as $x_{i}$ might have already satisfied the request. This can potentially make the algorithm run forever.
As mentioned earlier, coresets enable us to initialize the request sets $Q_{i}$ much more accurately due to the leader guessing phase. This, along with a crucial observation on the nature of unhappy pairs, allows us to bound the radii aspect ratio of requests in each $Q_{i}$ by $O(1/\varepsilon)$, whereas the loose upper bound approach of [ABB+23] bounds the radii aspect ratio only by $O(k/\varepsilon)$. This helps us in obtaining improved running time for uncapacitated $(k,z)$-clustering, as described above.
Finally, we note that a recent work of Huang et al. [HLLW25] gives a coreset framework for a number of constrained $(k,z)$-clustering objectives, including capacities and fairness, in several metric spaces. For continuous Euclidean, planar, and minor-free metrics, they design coresets for these objectives whose size is independent of $\log n$. In Section 4.4, we sketch an approach to obtain EPASes in the corresponding settings via essentially brute forcing on such coresets. However, this approach results in substantially worse running times as compared to our approach. Furthermore, although their framework designs coresets for many metric spaces that have bounded scatter dimension; it does not appear to give such a result for all metric spaces of bounded scatter dimension. For example, metrics of bounded highway dimension (which have bounded scatter dimension), do not seem to be captured by their framework. Therefore, in our opinion, their result along with brute force enumeration does not satisfactorily resolve 1.
2 Preliminaries
We consider metric (clustering) space $M=(P,F,{\sf d})$, where $P$ is a set of $n$ points/clients, and $F$ is a (possibly infinite) set of facilities, and ${\sf d}$ is a metric over $P\cup F$. Furthermore, a class $\mathcal{M}$ of metric spaces is a potentially infinite set of metric spaces. Note that, in case $F$ is infinite, the set $F$ may not be explicitly provided in the input.
For a real $B>1$ and $\delta>0$, we denote by $[B]_{\delta}$, the set containing $(1+\delta)^{j}$, for integers $0\leq j\leq\frac{2\log B}{\log(1+\delta)}$.
2.1 Scatter Dimension
In this section, we revise useful facts related to algorithmic scatter dimension. For a comprehensive background on scatter dimension, see [ABB+23].
Definition 7 (Ball Intersection Problem).
In the Ball intersection problem, we are given a metric space $M=(P,F,{\sf d})$ from a metric class $\mathcal{M}$, a finite set $Q\subsetneq P\times\mathbb{R}_{\geq 0}$ of point-distance constraints, called request set, and a an error parameter $\eta>0$. The goal is to find a center $x\in F$ such that ${\sf d}(p,x)\leq(1+\eta)\alpha$, for all $(p,\alpha)\in Q$, if such center exists, and report failure ()bottom($\bot$)( ⊥ ) otherwise.
We say $\mathcal{M}$ admits a ball intersection algorithm if there is an algorithm that correctly solves the ball intersection problem for every metric space $M\in\mathcal{M}$, and runs in time polynomial in the size [^9] When $F$ is infinite, then the size of $M$ is polynomial in $|P|$ and the space of storing a point. of $M$ and $1/\eta$.
Definition 8 (Algorithmic $\varepsilon$-Scatter Dimension).
[…]
  • $x_{i}=\texttt{BALL-INT}_{\mathcal{M}}(M,\{(p_{1},\alpha),\dots,(p_{i-1},\alpha_{i-1})\},\varepsilon/2)$$\forall 2\leq i\leq\kappa$
  • ${\sf d}(x_{i},p_{i})>(1+\varepsilon)\alpha_{i}$   $\forall i\in[\kappa]$
[…]
The following lemma is handy for bounding the number of iterations of our algorithm.
Lemma 9 [ABB+23, GI25]).
For any class $\mathcal{M}$ of metric space with algorithmic $\varepsilon$-scatter dimension $\lambda(\varepsilon)$, there exists a BALL-INT algorithm $\mathcal{C}_{\mathcal{M}}$ with the following property. Given $\varepsilon\in(0,1)$, a constant $t\geq 1$, and $a_{i}>0,\tau_{i}\geq 2$ for $i\in[t]$, any $(\mathcal{C}_{\mathcal{M}},\varepsilon)$-scattering contains $O(\sum_{i\in[t]}\lambda(\nicefrac{{\varepsilon}}{{2}})(\log\tau_{i})/\varepsilon)$ many triples whose radii lie in the interval $\cup_{i\in[t]}[a_{i},\tau_{i}a_{i}]$.
3 Framework
In this section, we present our EPAS framework for Constrained $(k,z)$-Clustering.
3.1 Formal Setup of the Framework.
In this paper, we consider $(k,z)$-clustering problems with assignment constraints, called Constrained $(k,z)$-Clustering. The general setup is as follows. Let $z\geq 1$ be a fixed integer. We are given an instance $\mathcal{I}$ of Constrained $(k,z)$-Clustering over a metric space $M=(P,F,{\sf d})$, and an integer $k\geq 1$. In addition, there may be problem-specific constraints (such as capacities, fairness, etc.). The goal is to select a set $X\subseteq F$ of size $k$ and an assignment $f:P\to X$, ensuring that both satisfy the given problem-specific constraints that could potentially depend on $X$. We refer to such $X$ as a feasible set of centers and such an $f$ as a feasible assignment from $P$ to $X$ The cost of pair $(X,f)$ — which we also refer to as a solution—is defined as ${\sf cost}(P,X,f)\coloneqq\sum_{p\in P}{\sf d}(p,f(p))^{z}$. Furthermore, the cost of a feasible set of $k$ centers $X\subseteq F$ is defined as ${\sf cost}(P,X)\coloneqq\min_{f}{\sf cost}(P,X,f)$, where the minimum is taken over all feasible assignments $f$. If no feasible assignment $f:P\to X$ exists, or if $X$ is not feasible, then we define ${\sf cost}(P,X)\coloneqq\infty$. The goal is to find a solution $(X,f)$ that minimizes the cost of the solution. In addition, our framework requires that the Constrained $(k,z)$-Clustering problem satisfies the following properties.
Coresets.
Consider a subset of clients $Y\subseteq P$, and a weight function $w:Y\to\mathbb{N}$. For a $k$-sized subset $C\subseteq F$ of centers, we say that $f:Y\times C\to\mathbb{N}$ is a feasible assignment from $Y$ to $C$ if (i) for each $p\in Y$, $\sum_{c\in C}f(p,c)=w(p)$, and (ii) $f$ satisfies additional problem-specific constraints [^10] Note that this definition of a feasible assignment naturally generalizes the definition of an assignment from an unweighted set $P$ to $X$, given earlier: if $f:P\to X$ is a feasible assignment, then construct an assignment $f^{\prime}:P\times X\to\left\{0,1\right\}$ by letting $f^{\prime}(p,c)=1$ iff $f(p)=c$. . We define the weighted-cost of $(Y,C,f)$ as ${\sf wcost}(Y,X,f)\coloneqq\sum_{p\in Y}\sum_{c\in C}f(p,c)\cdot{\sf d}(p,c)^{z}$. As before, we define the weighted-cost of $X$ as ${\sf wcost}(Y,C)\coloneqq\min_{f}{\sf wcost}(Y,C,f)$, where the min is taken over all feasible assignments $f:Y\times C\to\mathbb{N}$, and if no such feasible assignment exists, then ${\sf wcost}(Y,C,f)\coloneqq\infty$.
For an $0<\varepsilon\leq 1$, we say that $(Y,w)$ is an $\varepsilon$-coreset if, for each $C\subseteq F$ of size $k$, it holds that $|{\sf wcost}(Y,C)-{\sf cost}(P,C)|\leq\varepsilon\cdot{\sf cost}(P,C)$, with high probability.
We require that, for any $10<\varepsilon<$ 1, an $\varepsilon$-coreset for the Constrained $(k,z)$-clustering problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ is computable in polynomial-time in general metric spaces. We also allow the size of the coreset to depend on some problem-specific parameters (e.g., the number of colors in fair clustering), which will be reflected in the final running time.
Efficient computation of (near-) optimal assignments.
There exists an algorithm, ASSIGNMENT which takes as input a weighted point set $(Y,w)$, a candidate set $X\subseteq F$ of $k$ centers, and a parameter $10<\varepsilon<$ 1, and returns a feasible assignment $f:Y\times X\to\mathbb{N}$ such that ${\sf wcost}(Y,X,f)\leq(1+\varepsilon)\cdot{\sf wcost}(Y,X)$, where recall that ${\sf wcost}(Y,X)$ is the minimum-cost assignment from $Y$ to $X$. If no such assignment exists, then the algorithm returns $\bm{\bot}$. We require that the algorithm runs in time FPT in $k,\varepsilon$, and other problem-specific parameters. Note that, when this algorithm is used with the original (unweighted) point set $P$ with unit-weight function $\mathbbm{1}$, it returns a feasible assignment $f:P\times X\to\left\{0,1\right\}$, which can be interpreted as a feasible assignment $f:P\to X$ in a natural way. Thus, it suffices to compute a set of centers $X$, since a near-optimal assignment from either coreset $(Y,w)$ or the original point set $P$ to $X$ can be computed using ASSIGNMENT algorithm at the overhead of FPT time. We denote by $T_{\textsf{A}}(\mathcal{I})$ as the running time of ASSIGNMENT on input instance $\mathcal{I}$ [^11] For simplicity, we hide the dependence of the running time on $\varepsilon$.
(Feasible) Ball intersection algorithm.
[…]
We call such Constrained $(k,z)$-Clustering problem as Well Constrained $(k,z)$-Clustering problem. We allow each of the three algorithms to run in time FPT in $k,\varepsilon$, and potentially some problem-specific parameters. We use $|\mathcal{I}|$ to denote the size of the instance. [^12] When $F$ is infinite, then $|\mathcal{I}|$ is polynomial in input parameters other than $F$ and the space of storing a point.
3.2 Algorithm
Our main result is the following.
Theorem 10.
Let $\mathcal{M}$ be a class of metric spaces with algorithmic $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$. There is a EPAS that given an instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering over a metric space in $\mathcal{M}$ runs in time $2^{O\left(\tfrac{k}{\varepsilon}\log(k\ell)\cdot\lambda(\tfrac{\varepsilon}{60z})\log(\nicefrac{{1}}{{\varepsilon}})\right)}\operatorname{poly}(|\mathcal{I}|)\cdot T_{\textsf{A}}(\mathcal{I})\cdot T_{\textsf{B}}(\mathcal{I})$, where $\ell$ is the size of coreset for $\mathcal{I}$. Furthermore, the running time of the EPAS is $\exp{(O(\frac{k\lambda(\nicefrac{{\varepsilon}}{{60z}})}{\varepsilon})\log(\frac{1}{\varepsilon})\log(\frac{k\lambda(\nicefrac{{\varepsilon}}{{60z}})}{\varepsilon}))})\cdot|\mathcal{I}|^{O(1)}$ when $\ell=(\tfrac{k}{\varepsilon}\log|\mathcal{I}|)^{O(1)}$. Additionally, if coresets for $\mathcal{I}$ can be constructed deterministically, then the EPAS is deterministic.
The pseudo-code of our main algorithm yielding Theorem 10 is described in Algorithm 1. This algorithm is based on the ideas mentioned in Section 1.1, and it uses the recursive algorithm Algorithm 2 as a subroutine. It takes as input, an instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering, a ball intersection algorithm BALL-INT, an assignment algorithm ASSIGNMENT, and a guess $\mathcal{G}$ of optimal cost. Note that the guess, $\mathcal{G}$ of optimal cost, can be computed efficiently, up to a multiplicative factor of $(1+\varepsilon)$.
Algorithm 1 Approximation Scheme for Well Constrained $(k,z)$-Clustering
1:Input: Instance $\mathcal{I}$ of Well Constrained $(k,z)$-Clustering in metric space $M=(P,F,{\sf d})$, $\varepsilon\in(0,1/2]$, Ball intersection algorithm BALL-INT, ASSIGNMENT Algorithm and ${\sf OPT}\leq\mathcal{G}\leq(1+\varepsilon)\cdot{\sf OPT}$.
2:Output: A feasible solution $(S\subseteq F,f_{S})$ such that ${\sf cost}(P,S,f_{S})\leq(1+30\varepsilon){\sf OPT}$.
3:$(Y,w)\leftarrow$ coreset for $\mathcal{I}$
4:for every tuple $(p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{k})\in Y^{k}$ and distances $(r^{\prime}_{1},r^{\prime}_{2},\ldots,r^{\prime}_{k})\in[\Delta]_{\varepsilon}^{k}$ do
5:     Let $Q_{i}=\left\{(p^{\prime}_{i},r^{\prime}_{i})\right\}$ for all $i\in[k]$
6:     Let =$\mathcal{B}=\{B_{1},\dots,B_{k}\}$, where $B_{i}={\sf ball}(p^{\prime}_{i},\nicefrac{{6r^{\prime}_{i}}}{{\varepsilon}})$
7:     $S\leftarrow\texttt{COMPUTE-SOLUTION}((Q_{1},\dots,Q_{k}),\mathcal{B})$
8:     If $S\neq\bot$ then return $(S,\texttt{ASSIGNMENT}(P,S,\varepsilon))$
9:end for
10:return $\bot$
[…]
1:Let $X=(x_{1},\dots,x_{k})$ such that $x_{i}=\texttt{BALL-INT}(Q_{i},\varepsilon),i\in[k]$
2:return $\bot$ if no such $X$ was found
3:$f\leftarrow\texttt{ASSIGNMENT}(Y,X,(Q_{1},\dots,Q_{k}),\varepsilon)$, return $\bot$ if it fails
4:If ${\sf cost}(Y,X,f)\leq(1+5\varepsilon)\cdot\mathcal{G}$ then return $X$
5:for every index $i\in[k]$ and point $p\in B_{i}\setminus{\sf ball}(x_{i},r^{\prime}_{i})$ do
[…]
7:end for
8:return $\bot$
3.3 Analysis
In this section, we analyze the running time and correctness of Algorithm 1. First, in Section 3.3.1, we show that Algorithm 1 has bounded running time (Lemma 11), and later, in Section 3.3.2, we show that there there exists an iteration of for loop in 4 of Algorithm 1 that does not return $\bot$ (Lemma 12). Finally, we use these results to prove Theorem 10.
For keeping the analysis clean, we show the proof for Well Constrained $(k,z)$-Median. We remark that the same analysis can be extended easily to Well Constrained $(k,z)$-Clustering, by replacing $\varepsilon$ with $\varepsilon/z$. Let $\mathcal{I}$ be an instance of Well Constrained $(k,z)$-Clustering that is given as an input to Algorithm 1, and let $(Y,w)$ be the corresponding coreset for $\mathcal{I}$ obtained in Step 3.
3.3.1 Running time
In this section, we bound the running time of Algorithm 1.
Lemma 11.
Algorithm 1 terminates within time $2^{O(\tfrac{k}{\varepsilon}\log(k\ell)\cdot\lambda(\tfrac{\varepsilon}{2})\log(\nicefrac{{1}}{{\varepsilon}}))}\operatorname{poly}(|\mathcal{I}|)\cdot T_{\textsf{A}}(\mathcal{I})\cdot T_{\textsf{B}}(\mathcal{I})$.
[…]
Claim 11.1.
The depth of the recursion of Algorithm 1 is at most $O(k\lambda(\tfrac{\varepsilon}{2})\tfrac{\log(1/\varepsilon)}{\varepsilon})$.
Proof.
[…]
$\alpha^{j}_{i}=\frac{d(p^{j}_{i},x^{j}_{i})}{1+\varepsilon}\leq\frac{d(p^{j}_{i},p^{\prime}_{i})+d(p^{\prime}_{i},x^{j}_{i})}{1+\varepsilon}<\frac{6r^{\prime}_{i}}{\varepsilon}+d(p^{\prime}_{i},x^{j}_{i})\leq\frac{7r^{\prime}_{i}}{\varepsilon},$
[…]
Now, we bound the running time of Algorithm 1 as follows.
[…]
3.3.2 Correctness
In this section, we show that Algorithm 1 never returns $\bot$. Specifically, we show the following guarantee.
Lemma 12.
There exists an iteration of for loop (4) of Algorithm 1 whose COMPUTE-SOLUTION (7) never returns $\bot$.
Proof.
Let $\mathcal{O}=(O=(o_{1},\dots,o_{k}),f^{*})$ be an optimal solution to coreset $(Y,w)$ with cost ${\sf OPT}$. For $i\in[k]$, let $p^{*}_{i}=\arg\min_{p\in P}\{d(p,o_{i})\mid f^{*}\text{ assigns $p$ to $o_{i}$}\}$, be the leader of (the cluster corresponding to) $o_{i}$, by breaking ties arbitrarily. Similarly, let $r_{i}=d(p^{*}_{i},o_{i})$ be the distance of $o_{i}$ to its leader $p^{*}_{i}$. Now, we fix the iteration of the for loop of Algorithm 1 corresponding to $(p^{\prime}_{1},\dots,p^{\prime}_{k})=(p^{*}_{1},\dots,p^{*}_{k})$ and $(r^{\prime}_{1},\dots,r^{\prime}_{k})=(\tilde{r}_{1},\dots,\tilde{r}_{k})\in[\Delta]^{k}_{\varepsilon}$, where $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i},i\in[k]$. Let $\mathcal{T}$ be the recursion tree of COMPUTE-SOLUTION corresponding to this iteration with root $t_{0}$.
Observation 13.
At the root node $t_{0}$ of $\mathcal{T}$, we have that, for $i\in[k]$, $Q_{i}=(p^{*}_{i},\tilde{r}_{i})$, such that $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i}$.
We will show that there is at least one path $\mathcal{P}$ in $\mathcal{T}$ such that no nodes in $\mathcal{P}$ return $\bot$. Towards this, we define the following notion of consistency for a node in $\mathcal{T}$.
Definition 14 (Consistency).
A node $t\in\mathcal{T}$ with first argument $(Q_{1},\dots,Q_{k})$ is said to be consistent with $O$ if for any request $(p,\alpha)\in Q_{i},i\in[k]$, we have ${\sf d}(p,o_{i})\leq\alpha$.
Notice that if a node $t\in\mathcal{T}$ is consistent with $O$, then it successfully finds $X$ using BALL-INT in 1, since $o_{i}$ always satisfies all the requests in $Q_{i},i\in[k]$. Furthermore, since $f^{*}$ is feasible, ASSIGNMENT also finds a feasible assignment. Our aim, now, is to show that there is at least one path $\mathcal{P}$ in $\mathcal{T}$ such that all nodes in $\mathcal{P}$ are consistent. This means that none of the nodes in $\mathcal{P}$ return $\bot$, as desired.
Towards this, note that the root node $t_{0}\in\mathcal{T}$ has first argument as $((p^{\prime}_{1},\tilde{r}_{1}),\dots,(p^{\prime}_{k},\tilde{r}_{k}))$ and hence, is consistent with $O$, since for $(p^{\prime}_{i},\tilde{r}_{i})\in Q_{i}$, we have that ${\sf d}(p^{\prime}_{i},o_{i})=r_{i}\leq\tilde{r}_{i}$. Now, we show inductively that given a non-leaf node $t\in\mathcal{T}$ that is consistent with $O$, there exists at least one child $t^{\prime}$ of $t$ that is consistent with $O$. Let $(X,f)$ be the solution computed by node $t$ (such solution exists since $t$ is consistent with $O$). Then, the following claim is a key for consistency of $t^{\prime}$.
Claim 14.1.
If ${\sf cost}(Y,X,f)>(1+5\varepsilon)\cdot\mathcal{G}$, then there exists a point $p\in Y$ and an index $i\in[k]$ satisfying the following properties.
  1. (a)
    $f^{*}$ assigns $p$ to $i$,
  2. (b)
    ${\sf d}(p,x_{i})>(1+\varepsilon)\cdot{\sf d}(p,o_{i})$,
  3. (c)
    ${\sf d}(p,x_{i})\geq\tilde{r}_{i}$, and
  4. (d)
    ${\sf d}(p,p^{*}_{i})\leq\frac{6}{\varepsilon}\tilde{r}_{i}$.
Proof.
Let $W\coloneqq\left\{(p,i)\in Y\times[k]:\text{$f^{*}$ assigns $p$ to $i$, and }{\sf d}(p,x_{i})>(1+\varepsilon)\cdot{\sf d}(p,o_{i})\right\}$. First, we show that $W$ is non-empty. Suppose for the sake of contradiction that $U$ is empty, i.e., for each $p\in Y$ and all $i\in[k]$ such that $f^{*}$ assigns $p$ to $o_{i}$, it holds that ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$. Then, by the properties of BALL-INT and ASSIGNMENT algorithms, it follows that $(X,f^{*})$ is feasible. Hence, ${\sf cost}(Y,X,f^{*})\leq(1+\varepsilon){\sf cost}(Y,O,f^{*})$. Therefore, we have
${\sf cost}(Y,X,f)$ $\leq(1+\varepsilon)\cdot{\sf cost}(Y,X,f^{*})$ due to ASSIGNMENT
$\leq(1+\varepsilon)^{2}\cdot{\sf cost}(Y,O,f^{*})$
$\leq(1+\varepsilon)^{3}\cdot{\sf cost}(P,O,f^{*})$ due to coreset
$\leq(1+\varepsilon)^{3}\cdot\mathcal{G}$ since $\text{since }\mathcal{G}\geq{\sf OPT}$
$\leq(1+5\varepsilon)\cdot\mathcal{G}$ since $\text{since }\varepsilon\leq 1/2$
This contradicts the premise of the claim. Therefore, $W\neq\emptyset$.
Next, consider any pair $(p,i)\in W$, and note that, properties $(a)$ and $(b)$ are satisfied for $p$ and $i$, by definition. Next we show that $(p,i)$ also satisfy properties $(c)$ and $(d)$. For $(c)$, note that,
${\sf d}(p,x_{i})>(1+\varepsilon)\cdot{\sf d}(p,o_{i})=(1+\varepsilon)r_{i}>\tilde{r_{i}},$
as desired, since $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i}$ due to 13. Now, suppose the property $(d)$ does not hold, i.e., ${\sf d}(p,p^{*}_{i})>\frac{6\tilde{r}_{i}}{\varepsilon}$. Then, we obtain the following by triangle inequality.
${\sf d}(p,x_{i})\leq{\sf d}(p,p^{*}_{i})+{\sf d}(p^{*}_{i},x_{i})\leq{\sf d}(p,p^{*}_{i})+(1+\varepsilon)\tilde{r}_{i}\leq(1+\tfrac{\varepsilon}{4})\cdot{\sf d}(p,p^{*}_{i})$ (1)
Similarly, we obtain the following by triangle inequality.
${\sf d}(p,o_{i})\geq{\sf d}(p,p^{*}_{i})-{\sf d}(p^{*}_{i},o_{i})\geq{\sf d}(p,p^{*}_{i})-\tilde{r}_{i}\geq(1-\tfrac{\varepsilon}{6})\cdot{\sf d}(p,p^{*}_{i})$ (2)
Then, by combining equations (1) and (2), we obtain the following.
${\sf d}(p,x_{i})\leq\frac{(1+\frac{\varepsilon}{4})}{(1-\frac{\varepsilon}{6})}\cdot{\sf d}(p,o_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i}),$
contradicting the fact that $(p,i)\in W$. This shows that, for each pair $(p,i)\in W$, properties $(c)$ and $(d)$ hold. Since $W$ is non-empty, this completes the proof of the lemma. ∎
Fix some $(p_{t},i_{t})\in Y\times[k]$ satisfying 14.1 and note that $p_{t}\in B_{i_{t}}\setminus{\sf ball}(x_{i_{t}},\tilde{r}_{i_{t}})$. Hence, consider the corresponding iteration of the for loop to $(p_{t},i_{t})$ in 5, and let $t^{\prime}\in\mathcal{T}$ be the corresponding child of $t$ in this iteration. We claim that $t^{\prime}$ is consistent with $O$. Towards this, let $(Q_{1},\dots,Q_{k})$ and $(Q^{\prime}_{1},\dots,Q^{\prime}_{k})$ be first arguments at $t$ and $t^{\prime}$, respectively. Then, note that for $i\neq i_{t}$, we have, $Q^{\prime}_{i}=Q_{i}$, and $Q^{\prime}_{i_{t}}=Q_{i_{t}}\cup\{(p_{t},\alpha_{t})\}$, where $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}$. Hence, for any $i\neq i_{t}$, we have that for $(p,\alpha)\in Q^{\prime}_{i}=Q_{i}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Similarly, for $i_{t}$, we have for any $(p,\alpha)\in Q^{\prime}_{i_{t}}\setminus\{(p_{t},\alpha_{t})\}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Now, consider $(p_{t},\alpha_{t})\in Q^{\prime}_{i_{t}}$, and note that $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}>{\sf d}(p_{t},o_{i_{t}})$ since ${\sf d}(p_{t},x_{i_{t}})>(1+\varepsilon)\cdot{\sf d}(p_{t},o_{i_{t}})$ from 14.1, finishing the proof of the claim. Finally, note that since none of the nodes of path $\mathcal{P}$ return $\bot$, the root node $t_{0}$ never returns $\bot$, finishing the proof of the lemma.
Now, we are ready to finish the proof of Theorem 10.
Proof of Theorem 10.
Suppose Algorithm 1 terminates with a solution $(S,f_{S})$. Then, we have that ${\sf cost}(Y,S)\leq(1+5\varepsilon)\cdot\mathcal{G}\leq(1+5\varepsilon)(1+\varepsilon){\sf OPT}$. Therefore, we have
${\sf cost}(P,S,f_{S})$ $\leq(1+\varepsilon){\sf cost}(P,S)$ due to ASSIGNMENT
$\leq(1+\varepsilon)(1+2\varepsilon){\sf cost}(Y,S)$ due to coreset
$\leq(1+\varepsilon)^{2}(1+2\varepsilon)(1+5\varepsilon){\sf OPT}$
$\leq(1+30\varepsilon){\sf OPT},$
as desired.
Now, Lemma 11 says that Algorithm 1 terminates within time $2^{O\left(\tfrac{k}{\varepsilon}\log(k\ell)\cdot\lambda(\tfrac{\varepsilon}{60z})\log(\nicefrac{{1}}{{\varepsilon}})\right)}\operatorname{poly}(|\mathcal{I}|)\cdot T_{\textsf{A}}(\mathcal{I})\cdot T_{\textsf{B}}(\mathcal{I})$, and Lemma 12 implies that there is an iteration of for loop in 4 of Algorithm 1 that does not return $\bot$. Hence, Algorithm 1 terminates with a desired solution in time $2^{O\left(\tfrac{k}{\varepsilon}\log(k\ell)\cdot\lambda(\tfrac{\varepsilon}{60z})\log(\nicefrac{{1}}{{\varepsilon}})\right)}\operatorname{poly}(|\mathcal{I}|)\cdot T_{\textsf{A}}(\mathcal{I})\cdot T_{\textsf{B}}(\mathcal{I})$.
[…]
4 Applications of the Framework
In this section, we discuss some applications of the framework.
4.1 Capacitated Clustering
Uniform Capacities.
This is the most straightforward application of our framework. For capacitated $k$-median, a coreset of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ in [CL19a]. Note that this coreset is for general metric spaces, and even works with non-uniform capacities. Further, it is possible to generalize these coresets for $(k,z)$-clustering for any fixed $z\geq 1$, where the size of the coreset then depends on $z$. The Assignment subroutine in this case is simply the minimum-cost maximum flow problem, which is solvable in polynomial time, and produces integral flows. Finally, we note that uniform capacities is also the only sensible notion in continuous versions of capacitated $k$-median/means, where the set of candidate facilities is infinite, e.g., $\mathbb{R}^{d}$.
Non-Uniform Capacities.
Here, we need to slightly adapt the framework, for which we follow the approach taken in [CL19a]. We start with the coreset of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ from the same paper. Then, we use the technique of color coding, as follows. For each $c\in F$, we assign a color $i\in[k]$ chosen uniformly at random. A good coloring corresponds to a coloring in which all centers of a fixed optimal solution $O\subseteq F$ of size $k$ receieve distinct colors. Note that this happens with probability $1/e^{k}$, and this procedure can be derandomized by using standard tools from parameterized complexity [CFK+15]. Thus, now we assume that we have a good coloring, and we want to find a colorful set of $k$ centers. Then, in the algorithm, we make the following change. BALL-INT algorithm for the $i$th request set $Q_{i}$, selects a center that (i) satisfies all the requests in $Q_{i}$, and (ii) has a maximum capacity among all centers with color $i$. This ensures that, assuming all the guesses were correct, for each $i\in[k]$, we select a center $x_{i}$ such that (i) for each $p\in Y$, if $f^{*}$ assigns $p$ to $i$, then ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$, and (ii) capacity of $x_{i}$ is at least that of $o_{i}$. This guarantees that there is a feasible assignment from $Y$ to $X$ of cost at most $(1+O(\varepsilon))$ times ${\sf OPT}$.
4.2 $(\alpha,\beta)$-Fair Clustering and Variants
We begin with the definition of $(\alpha,\beta)$-fair $k$-median problem that has been studied in [CKLV17, BCJ+22, BFS24, HLLW25].
Definition 15.
The input consists of a metric space $\mathcal{M}=(P\cup F,d)$, where $P=P_{1}\cup P_{2}\cup\ldots P_{\ell}$, where $P_{i}$’s are (not necessarily disjoint) groups of clients in $P$. Further, the input also consists of two fairness vectors $\alpha,\beta\in[0,1]^{\ell}$. The objective is to select a set $X\subseteq F$ of $k$ centers, and an assignment $f:P\to X$ such that $f$ satisfies the following fairness constraints:
$\alpha_{i}\leq\frac{|\left\{x\in P_{i}:f(x)=c\right\}|}{|\left\{x\in P:f(x)=c\right\}|}\leq\beta_{i}\qquad\forall c\in X,\forall i\in[\ell]$
And further, ${\sf cost}(P,X)\coloneqq\sum_{p\in P}{\sf d}(p,f(p))$ is minimized among all such assignments. In this problem, $\Gamma$ denotes the number of equivalence classes of points w.r.t. groups, i.e., the number of distinct subsets of groups that a point can belong.
The following definitions pertaining to coresets for $(\alpha,\beta)$-fair clustering were introduced in [BFS24].
Definition 16 (Coloring Constraint and Universal Coresets, [BFS24]).
Fix an instance $(\mathcal{M}=(P\cup F,{\sf d}),(P_{1},P_{2},\ldots,P_{\ell}),\alpha,\beta,k)$ of the $(\alpha,\beta)$-fair $k$-median problem.
  • A coloring constraint is a $k\times\ell$ matrix $M$ containing non-negative entries.
  • Let $(Y,w)$ be a weighted set of points. Then, for a set of $k$ centers $X=\left\{c_{1},c_{2},\ldots,c_{k}\right\}$ and a coloring constraint $M\in\mathbb{N}^{k\times\ell}$, ${\sf wcost}(Y,X,M)$ is defined as the minimum value of $\sum_{p\in Y}\sum_{i\in[k]}f(p,c_{i})\cdot{\sf d}(p,c_{i})$, over all assignments $f:Y\times X\to\mathbb{N}$ satisfying the following constraints.
    • For each $p\in P$, $\sum_{i\in[k]}f(p,c_{i})=w(p)$, and
    • For each $i\in[k]$ and each group $j\in[\ell]$, $\sum_{p\in P_{j}}f(p,c_{i})=M_{ij}$.
    If no such assignment exists, then ${\sf wcost}(Y,X,M)$ is defined as $\infty$. Further, when $(Y,w)$ is the original set of points $P$ with unit weight function $\mathbbm{1}$, then ${\sf wcost}(P,X,M)$ is written as ${\sf cost}(P,X,M)$.
  • A universal coreset is a pair $(Y,w)$, such that for every set $X$ of $k$ centers and every coloring constraint $M$, the following holds:
    $(1-\varepsilon)\cdot{\sf cost}(P,X,M)\leq{\sf wcost}(Y,X,M)\leq(1+\varepsilon)\cdot{\sf cost}(P,X,M).$
Universal coresets of size $\Gamma\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ in general metrics were designed in [BFS24], and the extensions to $(k,z)$-clustering variant (which is defined by naturally modifying the definitions above by including the $z$-th power of distances) with improved size were given in [HLLW25]. For our purpose, the former guarantee on the size suffices. Further, [BFS24] also designed an algorithm that takes input a weighted point set $(Y,w)$, and a set $X=\left\{c_{1},c_{2},\ldots,c_{k}\right\}$ of $k$ centers, runs in time $(k\Gamma)^{O(k\Gamma)}\cdot n^{O(1)}$, and returns an an assignment $f:Y\times X\to\mathbb{N}$, such that the cost $\sum_{p\in Y}\sum_{i\in[k]}f(p,c_{i})\cdot{\sf d}(p,c_{i})$ is minimized across all assignments that respect the $\alpha,\beta$ fairness constraints as above. Further, it is easy to check that this algorithm does not depend on the objective function – in fact it can compute an optimal assignment for any $(k,z)$-clustering objective. Thus, we have all the ingredients to plug into our framework for $(\alpha,\beta)$-fair $(k,z)$-clustering in metric spaces of bounded scatter dimension for any fixed $z\geq 1$, and obtain an EPAS with running time $\Gamma^{O(\Gamma k\lambda(\varepsilon)\log(\frac{1}{\varepsilon})(\log k+\log(\lambda(\varepsilon))))}\cdot(|\mathcal{I}|)^{O(1)}$.
Further implications.
[BFS24] discuss the following variants of constrained clustering in their paper.
Lower Bounded Clustering.
Here, the points have only a single color, but the size of each cluster must be at least $L$, where $L$ is a non-negative integer given in the input.
$\ell$-Diversity Clustering.
Here, the points are again classified in one of $\ell$ groups, $P_{1},P_{2},\ldots,P_{\ell}$ (a point can belong to multiple groups), and each cluster $A$ should satisfy that $|A\cap P_{\ell}|\leq\frac{|A|}{\ell}$.
Chromatic Clustering.
The setting is similar to above, but each cluster can contain at most one point of each group.
[BFS24] discuss in detail the modifications (if any) to the universal coreset and the FPT algorithm for finding an optimal assignment required to handle each of these variants, and we direct the reader to these discussions. Since these are the only two ingredients required in our problem, our framework implies an EPAS for all of these variants in metric spaces of bounded scatter dimension.
4.3 Vanilla (Voronoi) Clustering
Our framework also naturally for vanilla clustering, implying an EPAS with running time $2^{O_{\varepsilon}(k\log k)}\cdot(|\mathcal{I}|)^{O(1)}$. Notice that this already improves upon the $2^{O_{\varepsilon}(k\log^{2}k)}\cdot n^{O(1)}$ running time from [ABB+23], where $O_{\varepsilon}(\cdot)$ notation hides the dependence on $\varepsilon$. However, the dependence on $\varepsilon$ may be worse for metric spaces with super-exponential scattter dimension. Here, we describe a simple modification for choosing a point of the coreset in the algorithm (as opposed to the for loop of line 4 in Algorithm 2) that, (i) has improved dependence on $k$ as described above, and (ii) matches the dependence on $\varepsilon$ as in [ABB+23], thus improving the running time as compared to [ABB+23] in all regimes.
To this end, first we make the following observation: in Voronoi clustering, without loss of generality, it can be assumed that for each $p\in Y$, the entire weight $w(p)$ is assigned to a single cluster-center $x_{i}\in X$ that is closest to $p$. Then, in Algorithm 2 instead of the for loop of lines 4 to 7: we sample a point $p\in\bigcup_{i\in[k]}B_{i}$ according to the following distribution: $\Pr(p=a)=\frac{w(a){\sf d}(a,X)}{\sum_{p\in\cup_{i}B_{i}}w(p){\sf d}(p,X)}$ for $a\in\cup_{i\in[k]}B_{i}$. Let $p$ be the sampled point. Next, we also sample an index $i$ uniformly at random from the set $\left\{j\in[k]:p\in B_{i}\right\}$. Then, we make a similar recursive call to the same algorithm by adding the request $(p,\frac{{\sf d}(p,x_{i})}{1+\varepsilon/3})$ to $Q_{i}$, similar to line 5. The crucial observation is the following, which is inspired from similar claims from [ABB+23, GI25].
Claim 16.1.
Let $X$ be a solution such that ${\sf cost}(Y,X)>(1+5\varepsilon)\cdot\mathcal{G}$, and let $W\coloneqq\{p\in Y:{\sf d}(p,X)>(1+\varepsilon)\cdot{\sf d}(p,O)\}$. Then, (i) $W\subseteq\bigcup_{i\in[k]}B_{i}$, and (ii) $\frac{\sum_{p\in W}w(p){\sf d}(p,X)}{\sum_{p\in Y}w(p){\sf d}(p,X)}\geq\frac{\varepsilon}{10}$.
Proof.
Consider $X$ and the corresponding set $W$ as in the statement. Let $H\coloneqq Y\setminus W$, and note that for all points $p\in H$, ${\sf d}(p,X)\leq(1+\varepsilon)\cdot{\sf d}(p,O)$. For any $T\subseteq Y$, let $C_{T}\coloneqq\sum_{p\in T}w(p){\sf d}(p,X)$.
First, let us suppose that there exists a point $p\in W$ such that ${\sf d}(p,x_{i})>\frac{6r^{\prime}_{i}}{\varepsilon}$ for all $i\in[k]$. Then, by arguments similar to that in 14.1, we can argue that ${\sf d}(p,X)<(1+\varepsilon){\sf d}(p,O)$, contradicting the assumption that $p\in W$. This shows part (i).
Let us now suppose for contradiction for part (ii) that $\frac{C_{W}}{C_{Y}}<\frac{\varepsilon}{10}$.Then, consider:
$C_{H}=\sum_{p\in H}w(p){\sf d}(p,X)$ $\leq(1+\varepsilon)\cdot\sum_{p\in H}w(p){\sf d}(p,O)\leq(1+\varepsilon){\sf OPT}$
Now, note that $C_{Y}=C_{H}+C_{W}\leq(1+\varepsilon){\sf OPT}+\frac{\varepsilon}{10}C_{Y}$, which implies that $(1-\frac{\varepsilon}{10})C_{Y}\leq(1+\varepsilon){\sf OPT}\implies C_{Y}\leq\frac{1+\varepsilon}{1-\varepsilon/10}{\sf OPT}\leq(1+2\varepsilon){\sf OPT}\leq(1+2\varepsilon)\mathcal{G}$, contradicting the hypothesis. ∎
Assuming this claim, we know that the probability that the point $p$ sampled according to the aforementioned distribution belongs to $W$ is at least $\frac{\varepsilon}{10}$. Conditioned on this event, with probability at least $\frac{1}{k}$, the sampled cluster index $i$ in the aforementioned way is same as that of $p$ in the optimal solution $O$. Hence, the state of the algorithm remains consistent with $O$. The depth of the recursion remains bounded by $O(k\log(\frac{1}{\varepsilon})\cdot\lambda(\varepsilon))$, which implies that we obtain an algorithm with running time $2^{O(k\log(\frac{1}{\varepsilon})\lambda(\varepsilon/c)\cdot\log(\frac{k}{\varepsilon}))}\cdot(|\mathcal{I}|)^{O(1)}$, for some constant $c$ that depends only on $z$.
4.4 Additional Remarks
Handling Outliers.
Jaiswal and Kumar [JK23] (also [DGI25]), extending the work of Agrawal et al. [AISX23] gave a general additive-$\varepsilon$ approximation-preserving reduction from constrained $(k,z)$-clustering with $m$ outliers to its analogous constrained version without outliers that has a multiplicative overhead of $\left(\frac{k+m}{\varepsilon}\right)^{O(m)}\cdot n^{O(1)}$. By plugging in our EPASes for constrained clustering into this result, we also obtain EPASes for the corresponding outlier versions.
Relations to Improved Coresets for Constrained Clustering.
For most known metric spaces with bounded scatter dimension, the recent work of Huang et al. [HLLW25] gives coresets of size $\text{poly}(k,\varepsilon)$ for a large class of constrained $(k,z)$-clustering problems, including capacitated clustering, $(\alpha,\beta)$-fair clustering, fault-tolerant clustering, and more. To put this work into a proper context w.r.t. our EPASes, two remarks are in order.
Firstly, we note that such coresets can be used to obtain EPASes for the corresponding problems [^13] Although the following argument is very simple and should be folklore, we are unable to find a reference that explicitly describes how coresets of size independent of $\log n$ may be used to obtain EPASes for constrained $(k,z)$-clustering. Therefore, we provide the argument for the sake of completeness in order to perform a fair comparison with our results.. Let $(Y,w)$ be the coreset of size $\text{poly}(k,\varepsilon)$. Then, for each $p\in Y$ and each $i\in[k]$, we guess (up to a factor of $(1+\varepsilon)$) the portion of weight of $p$ assigned to the $i$th cluster. Note that the number of guesses is bounded by $(\frac{\log n}{\varepsilon})^{k\cdot|Y|}$, which is FPT in $k$ and $\varepsilon$. Fix one such guess. Then, for each cluster $i\in[k]$, we can find a feasible center that optimizes the cost (up to a factor of $(1+\varepsilon)$ arising from the guessing). Let $X$ be the set of centers chosen in this manner. We compute an optimal assignment using the assignment algorithm in time FPT in $k$ and $\varepsilon$. Finally, we return the minimum-cost solution found over all guesses. Although this strategy also yields an EPAS for many of the constrained problems in metric spaces of bounded scatter dimension, we note that the running time is $(k|Y|/\varepsilon)^{O(k|Y|)}\cdot n^{O(1)}$, and the size of coresets is $\Omega_{\varepsilon}(k^{2})$. Therefore, the dependence on $k$ of the running time is $2^{\Omega_{\varepsilon}(k^{3}\log k)}\cdot n^{O(1)}$, which is much worse compared to our running time of $2^{O_{\varepsilon}(k\log k)}\cdot n^{O(1)}$. Secondly, although the result of [HLLW25] gives coresets of size independent of $\log n$ in most known metric spaces of bounded scatter dimension, it is not known to yield such coresets for all metric spaces of bounded scatter dimension. Thus, this result does not answer the open question of [ABB+23] in its entirety.
Secondly, we note that plugging in the improved coresets into our algorithm improves the dependence on $\varepsilon$ in running time of the algorithm – a factor of $O(\log(\lambda(\varepsilon)))$ in the exponent is replaced by a factor of $O(\log(\frac{1}{\varepsilon}))$. Since the scatter dimension $\lambda(\varepsilon)$ of all of these classes of metric spaces is bounded by $2^{1/\varepsilon^{O(1)}}$, this results in an improved dependence on $\varepsilon$.
Acknowledgments.
We thank Andreas Emil Feldmann for bringing to our attention the work [FV22], which rules out the existence of an EPAS for the Capacitated $k$-Center problem in metrics of bounded highway dimension.